40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Dec 17 01:56:16 2011, Koji, Summary, SUS, Hysteresis test 
    Reply  Sat Dec 17 03:59:32 2011, kiwamu, Update, SUS, Aborted Hysteresis test 
       Reply  Sat Dec 17 11:53:46 2011, Zach, Update, SUS, Aborted Hysteresis test 
          Reply  Sat Dec 17 12:41:46 2011, Koji, Update, SUS, Aborted Hysteresis test 
       Reply  Sun Dec 18 16:16:55 2011, kiwamu, Update, SUS, Another trial of Hysteresis test 
          Reply  Sun Dec 18 19:56:00 2011, kiwamu, Update, SUS, Another trial of Hysteresis test 
Message ID: 6129     Entry time: Sat Dec 17 03:59:32 2011     In reply to: 6128     Reply to this: 6130   6132
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: Aborted Hysteresis test 

Quote from #6128

To test it, we are shaking all of the suspension biases +/-1.0 with a script.

The hysteresis test has been aborted.

All of the suspensions have accumulated unexpectedly big DC biases of about 5 from their nominal points.

In fact the ITMX and ITMY mirrors started being stacked to their OSEMs.
The script process has been force-quit and I have restored all the DC biases to their nominal points.
They still look okay: MC can be locked at the 00 mode, DRMI fringe is visible at AS, the green beams are resonating the arm cavities
Need another trial.
ELOG V3.1.3-